# MaxSup: Overcoming Representation Collapse in Label Smoothing

### v2

@@ 4.1 Analysis of MaxSup’s Learning Benefits
Having established how MaxSup addresses Label Smoothing’s (LS) principal shortcomings (Section 3.1), we now demonstrate its impact on inter-class separability and intra-class variation—two properties essential for accurate classification and effective transfer learning.
Table 2: Feature quality metrics for ResNet-50 on ImageNet-1K. While all methods reduce $\bar{d}_{\text{within}}$ (intra-class variation) relative to the baseline, MaxSup retains the greatest within-class diversity. Both $\bar{d}_{\text{within}}$ and $R^{2}$ benefit from higher values.
@@ 4.1.1 Intra-Class Variation and Transferability
Table 2 compares intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$[13] for ResNet-50 trained on ImageNet-1K. Although all investigated regularizers decrease $\bar{d}_{\text{within}}$ relative to a baseline, MaxSup yields the smallest reduction, indicating a stronger retention of subtle within-class diversity—widely associated with enhanced generalization and improved transfer performance.
These benefits are further underscored by the linear-probe transfer accuracy on CIFAR-10 (Table 3). While LS and Logit Penalty each boost ImageNet accuracy, both degrade transfer accuracy, likely by suppressing informative and transferable features. By contrast, MaxSup preserves near-baseline performance, implying that it maintains rich discriminative information crucial for downstream tasks. For extended evaluations on diverse datasets, see Table 9 in the appendix.
@@ 4.1.2 Connection to Logit Penalty
As detailed in Section 3, both Label Smoothing (LS) variants and MaxSup impose penalties directly at the logit level, aligning with the perspective that various regularizers influence a model’s representational capacity via distinct logit constraints [13]. Within this family of techniques, Logit Penalty and MaxSup both address the maximum logit, yet diverge fundamentally in their specific methods of regularization.
Logit Penalty minimizes the $\ell_{2}$-norm of the entire logit vector, inducing a global contraction that can improve class separation but also reduce intra-class diversity, potentially hindering downstream transfer. By contrast, MaxSup focuses exclusively on the top-1 logit, gently nudging it closer to the mean logit. Because only the highest-confidence prediction is penalized, MaxSup avoids the uniform shrinkage observed in Logit Penalty, preserving richer intra-class variation—a property essential for robust transfer. Further insights into this behavior can be found in Appendix N, where logit-value histograms illustrate how each method affects the logit distribution.
@@ 4.2.1 Experiment Setup
Model Training Configurations.We evaluate both convolutional (ResNet[11], MobileNetV2 [25]) and transformer (DeiT-Small [33]) architectures on ImageNet [15]. For the ResNet Series, we train for 200 epochs using stochastic gradient descent (SGD) with momentum0.9, weight decay of $1\times 10^{-4}$, and a batch size of 2048. The initial learning rate is 0.85 and is annealed via a cosine schedule. [^2] Additional training hyperparameters follow the FFCV scripts at https://github.com/libffcv/ffcv. See Appendix E for further details. We also test ResNet variants on CIFAR-100 with a conventional setup: an initial learning rate of 0.1 (reduced fivefold at epochs 60, 120, and 160), training for 200 epochs with batch size 128 and weight decay $5\times 10^{-4}$. For DeiT-Small, we use the official codebase [33], training from scratch without knowledge distillation to isolate MaxSup’s contribution. CutMix and Mixup are disabled to ensure the model optimization objective remains unchanged.
Hyperparameters for Compared Methods.We compare Max Suppression Regularization against a range of LS extensions, including Zipf Label Smoothing[19] and Online Label Smoothing [41]. Where official implementations exist, we adopt them directly; otherwise, we follow the methodological details provided in each respective paper. Except for any method-specific hyperparameters, all other core training settings remain identical to the baselines. Furthermore, both MaxSup and standard LS employ a linearly increasing $\alpha$-scheduler for improved training stability (see Appendix F). This ensures a fair comparison under consistent and reproducible training protocols.
@@ 4.2.2 Experiment Results
ConvNet Comparison. Table4 shows results for MaxSup alongside various label-smoothing and self-distillation methods on both ImageNet and CIFAR-100 benchmarks. Across all convolutional architectures tested, MaxSup consistently delivers the highest top-1 accuracy among label-smoothing approaches. By contrast, OLS [41] and Zipf-LS [19] exhibit less stable gains, suggesting their effectiveness may heavily hinge on specific training protocols.
@@
Label Smoothing 76.08 ($+1.69$) 0.16
Zipf-LS 75.89 ($+1.50$) 0.26
OLS 76.16 ($+1.77$) 0.18
MaxSup 76.49 ($+2.10$) 0.12
@@
Fine-Grained Classification. Beyond ImageNet, we evaluate MaxSup on demanding fine-grained benchmarks such as CUB-200 [34] and Stanford Cars [14]. As shown in Table 11, MaxSup outperforms other LS-based methods by more effectively capturing subtle within-class variations, which are critical for success in fine-grained recognition scenarios.
Ablation on the Weight Schedule. We also systematically investigate how different $\alpha$ scheduling strategies impact MaxSup’s performance. Empirical results indicate that MaxSup consistently maintains high accuracy across a wide range of schedules, further underscoring its robustness against hyperparameter changes. For additional details and discussions, refer to Appendix F.
@@
Label Smoothing 42.4 ($+0.3$)
MaxSup 42.8 ($+0.7$)
@@ 4.3 Evaluation on Semantic Segmentation
We further investigate MaxSup’s applicability to downstream tasks by evaluating its performance on semantic segmentation using the widely adopted MMSegmentation framework. [^3] https://github.com/open-mmlab/mmsegmentation Specifically, we adopt the UperNet [36] architecture with a DeiT-Small backbone, trained on ADE20K. Models pretrained on ImageNet-1K with either MaxSup or Label Smoothing are then fine-tuned under the same cross-entropy objective (Section 4.2.2).
@@
Figure 2: Grad-CAM [27] visualizations for DeiT-Small models under three training setups: MaxSup (2nd row), Label Smoothing (3rd row), and a baseline (4th row). The first row shows the original images. Compared to Label Smoothing, MaxSup more effectively filters out non-target regions and highlights essential features of the target class, reducing instances where the model partially or entirely focuses on irrelevant areas.
@@ 4.4 Visualization via Class Activation Maps
To better understand how MaxSup fundamentally differs from Label Smoothing (LS) in guiding model decisions, we employ Gradient-weighted Class Activation Mapping (Grad-CAM) [27], which highlights regions most influential for each prediction.
We evaluate DeiT-Small under three training setups: MaxSup (second row), LS (third row), and a baseline with standard cross-entropy (fourth row). As illustrated in Figure 2, MaxSup-trained models more effectively suppress background distractions than LS, which often fixates on unrelated objects—such as poles in “Bird,” tubes in “Goldfish,” and caps in “House Finch.” This behavior reflects LS’s error-enhancement mechanism, which can misdirect attention.
Moreover, MaxSup retains a wider spectrum of salient features, as exemplified in the Shark” and Monkey” images, where LS-trained models often omit crucial semantic details (e.g., fins, tails, or facial contours). These findings align with our analysis in Appendix J, clearly demonstrating that MaxSup preserves richer intra-class information. Consequently, MaxSup-trained models produce more accurate and consistent predictions by effectively leveraging fine-grained object cues. Further quantitative Grad-CAM overlay metrics (e.g., precision and recall for target regions) confirm that MaxSup yields more focused and comprehensive activation maps, further underscoring its overall efficacy.
@@ 5 Conclusion
Limitations. Although promising, several questions remain. Prior work[21] notes that LS-trained teachers may degrade knowledge distillation [12], and Guo et al. [9] suggests LS accelerates convergence via conditioning. Examining MaxSup’s role in distillation and its impact on training dynamics would clarify these effects. Recent studies [31, 7] also show that $\ell_{2}$ regularization biases final-layer features toward low-rank solutions, raising questions about whether MaxSup behaves similarly.
@@ 63. Answer: [Yes]
Appendix A Proof of Lemma 3.2
@@
Appendix B Proof of Theorem 3.3
@@
Appendix C Gradient Analysis
C.1 New Objective Function
@@ C.2 Gradient Analysis
Appendix D Pseudo Code
@@
Appendix E Robustness Under Different Training Recipes
@@
Appendix F Increasing Smoothing Weight Schedule
@@
As shown in Table 8, both Label Smoothing and MaxSup benefit from this $\alpha$ scheduler. For Label Smoothing, accuracy improves from 75.91% to 76.16%, while MaxSup sees a more pronounced gain, from 76.12% to 76.58%. This greater improvement for MaxSup ($+0.46\%$) compared to Label Smoothing ($+0.25\%$) corroborates our claim that MaxSup successfully addresses the inconsistent regularization and error-enhancement issues of Label Smoothing during misclassifications.
@@
Appendix G Extended Evaluation of Linear Transferability on Different Datasets
To further demonstrate the substantial improvement in feature representation compared to other methods, we further compare the linear transfer accuracies of different methods on a broader range of datasets in Table 9
Table 9: Validation performance of different methods, evaluated using multinomial logistic regression with l2 regularization. Although Label Smoothing and OLS improve ImageNet accuracy, they substantially degrade transfer accuracy compared to MaxSup. Following [13], we selected from 45 logarithmically spaced values between $10^{-6}$ and $10^{5}$. Notably, the search range is larger than the search range used in Table 3, thus leading to higher overall accuracies on CIFAR10.
@@ Appendix G Extended Evaluation of Linear Transferability on Different Datasets
Appendix H Extended Comparison to More Label Smoothing Alternatives
We have included Confidence Penalty [23] and Adaptive Label Smooothing with Self-Knowledge [18] for an extended comparison in Table 10. We follow their recommended hyperparameter settings: weight coefficient is set to 0.1 for confidence penalty, and the checkpoint with the highest validation accuracy checkpoint is treated as teacher for Adaptive Label Smoothing. From Table 10, we see that MaxSup outperforms both Confidence Penalty and Adaptive LS on ImageNet with ResNet-50. This aligns with our theoretical analysis that selectively penalizing $z_{\mathrm{max}}$ yields a more consistent regularization than penalizing all logits (Confidence Penalty) or dynamically smoothing the label distribution with self knowledge (Adaptive LS).
Table 10: Comparison of classic convolutional neural networks on ImageNet and CIFAR-100. Results are reported as “mean ± std” (percentage). Bold entries highlight the best performance; underlined entries mark the second best. (Methods with denote code adaptations from official repositories; see text for details.)

### v3

@@ 4.1 Analysis of MaxSup’s Learning Benefits
Having established how MaxSup addresses Label Smoothing’s (LS) principal shortcomings (Section˜3.1), we now demonstrate its impact on inter-class separability and intra-class variation—two properties essential for accurate classification and effective transfer learning.
Table 2: Feature quality of ResNet-50 on ImageNet-1K.
@@ 4.1.1 Intra-Class Variation and Transferability
Table 2 compares intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$[kornblith2021better] for ResNet-50 trained on ImageNet-1K. Although all investigated regularizers decrease $\bar{d}_{\text{within}}$ relative to a baseline, MaxSup yields the smallest reduction, indicating a stronger retention of subtle within-class diversity—widely associated with enhanced generalization and improved transfer performance.
These benefits are further underscored by the linear-probe transfer accuracy on CIFAR-10 (Table 3). While LS and Logit Penalty each boost ImageNet accuracy, both degrade transfer accuracy, likely by suppressing informative and transferable features. By contrast, MaxSup preserves near-baseline performance, implying that it maintains rich discriminative information crucial for downstream tasks. For extended evaluations on diverse datasets, see Table˜12 in the appendix.
@@ 4.1.2 Connection to Logit Penalty
As detailed in Section 3, both Label Smoothing (LS) variants and MaxSup impose penalties directly at the logit level, aligning with the perspective that various regularizers influence a model’s representational capacity via distinct logit constraints [kornblith2021better]. Within this family of techniques, Logit Penalty and MaxSup both address the maximum logit, yet diverge fundamentally in their specific methods of regularization.
Logit Penalty minimizes the $\ell_{2}$-norm of the entire logit vector, inducing a global contraction that can improve class separation but also reduce intra-class diversity, potentially hindering downstream transfer. By contrast, MaxSup focuses exclusively on the top-1 logit, gently nudging it closer to the mean logit. Because only the highest-confidence prediction is penalized, MaxSup avoids the uniform shrinkage observed in Logit Penalty, preserving richer intra-class variation—a property essential for robust transfer. Further insights into this behavior can be found in Appendix˜M, where logit-value histograms illustrate how each method affects the logit distribution.
@@ 4.2.1 Experiment Setup
Model Training Configurations.We evaluate both convolutional (ResNet[he2016deep], MobileNetV2 [sandler2018mobilenetv2]) and transformer (DeiT-Small [touvron2021training]) architectures on ImageNet [krizhevsky2012imagenet]. For the ResNet Series, we train for 200 epochs using stochastic gradient descent (SGD) with momentum0.9, weight decay of $1\times 10^{-4}$, and a batch size of 2048. The initial learning rate is 0.85 and is annealed via a cosine schedule. [^2] Additional training hyperparameters follow the FFCV scripts at https://github.com/libffcv/ffcv. See Appendix F for further details. We also test ResNet variants on CIFAR-100 with a conventional setup: an initial learning rate of 0.1 (reduced fivefold at epochs 60, 120, and 160), training for 200 epochs with batch size 128 and weight decay $5\times 10^{-4}$. For DeiT-Small, we use the official codebase [touvron2021training], training from scratch without knowledge distillation to isolate MaxSup’s contribution. CutMix and Mixup are disabled to ensure the model optimization objective remains unchanged.
Hyperparameters for Compared Methods.We compare Max Suppression Regularization against a range of LS extensions, including Zipf Label Smoothing[liang2022efficient] and Online Label Smoothing [zhang2021delving]. Where official implementations exist, we adopt them directly; otherwise, we follow the methodological details provided in each respective paper. Except for any method-specific hyperparameters, all other core training settings remain identical to the baselines. Furthermore, both MaxSup and standard LS employ a linearly increasing $\alpha$-scheduler for improved training stability (see Appendix˜G). This ensures a fair comparison under consistent and reproducible training protocols.
@@ 4.2.2 Experiment Results
ConvNet Comparison. Table4 shows results for MaxSup alongside various label-smoothing and self-distillation methods on both ImageNet and CIFAR-100 benchmarks. Across all convolutional architectures tested, MaxSup consistently delivers the highest top-1 accuracy among label-smoothing approaches. By contrast, OLS [zhang2021delving] and Zipf-LS [liang2022efficient] exhibit less stable gains, suggesting their effectiveness may heavily hinge on specific training protocols.
@@
Label Smoothing 76.08 (+1.69+1.69) 0.16
Zipf-LS 75.89 (+1.50+1.50) 0.26
OLS 76.16 (+1.77+1.77) 0.18
MaxSup 76.49 (+2.10+2.10) 0.12
@@
Table 6: Classification on CUB and Cars Datasets.
Method CUB[cub] Cars[car]
Baseline 80.88 90.27
LS 81.96 91.64
OLS 82.33 91.96
Zipf-LS 81.40 90.99
MaxSup 82.53 92.25
Fine-Grained Classification. Beyond large-scale benchmarks like ImageNet, we further evaluate MaxSup on two fine-grained visual recognition tasks: CUB-200-2011 [cub] and Stanford Cars [car]. These datasets pose unique challenges due to subtle inter-class differences, which often expose the limitations of standard regularization approaches. As shown in Table˜6, MaxSup achieves the best performance across both datasets, surpassing LS and its recent variants. This demonstrates that MaxSup encourages the model to learn more discriminative and semantically rich representations that better capture fine-grained attributes, such as textures and part-level details. The consistent improvements on these benchmarks further validate MaxSup’s capacity to generalize across different visual domains and its potential to enhance robustness in recognition scenarios where nuanced feature understanding is critical.
Table 7: Comparison of classification performance (%) across imbalance levels for different loss strategies (Focal Loss vs Label Smoothing (LS) vs MaxSup) on the long-tailed CIFAR-10 dataset using Resnet-32. Best performances are in bold.
Dataset Split Imbalance Ratio Method Overall Many Medium Low
LT CIFAR-10 val 50 Focal Loss 77.4 76.0 89.7 0.0
Label Smoothing 81.2 81.6 77.0 0.0
MaxSup 82.1 82.5 78.1 0.0
LT CIFAR-10 test 50 Focal Loss 76.8 75.3 90.4 0.0
Label Smoothing 80.5 81.1 75.4 0.0
MaxSup 81.4 82.3 73.4 0.0
LT CIFAR-10 val 100 Focal Loss 75.1 71.8 88.3 0.0
Label Smoothing 76.6 80.6 60.7 0.0
MaxSup 77.1 80.1 65.1 0.0
LT CIFAR-10 test 100 Focal Loss 74.7 71.6 87.2 0.0
Label Smoothing 76.4 80.8 59.0 0.0
MaxSup 76.4 79.9 62.4 0.0
Long-Tailed Classification. To assess the effectiveness of MaxSup under data imbalance, we performed experiments on the CIFAR-10-LT dataset with imbalance ratios of 50 and 100, following the experimental settings described in [tang2020long]. The corresponding results are summarized in Table˜7. The evaluation compares three setups: Focal Loss, Focal Loss + LS, and Focal Loss + MaxSup. Across all imbalance ratios and splits (val/test), MaxSup consistently outperforms both the baseline and LS in overall accuracy, which jointly reflects the many-shot, medium-shot, and low-shot (minor class) performance. For example, at an imbalance ratio of 50 on the test split, MaxSup achieves $81.4\%$ accuracy, outperforming Focal Loss ($76.8\%$) by 4.6 percentage points, and LS ($80.5\%$) by 0.9 percentage points. These results indicate that MaxSup achieves a better trade-off between many- and medium-shot accuracy. While it does not fully resolve the challenge of imbalanced classification—especially for minority classes—it shows positive effects and offers a promising direction for further extension.
Table 8: Comparison of MaxSup, Label Smoothing (LS), and standard Cross Entropy (CE) on CIFAR-10-C. Lower is better. Values show mean(std) across three setups.
Metric MaxSup LS CE
Error (Corr) 0.362(0.055) 0.359(0.064) 0.354(0.015)
NLL (Corr) 1.770(0.103) 1.476(0.111) 1.819(0.158)
ECE (Corr) 0.145(0.003) 0.158(0.015) 0.260(0.015)
Corrupted Image Classification To evaluate the effectiveness of MaxSup on out-of-distribution (OOD) settings, we also conducted experiments on CIFAR10-C benchmark [hendrycksbenchmarking] shown in Table˜8 following settings in [heinonen2025robust]. Table˜8 reports the performance of MaxSup and Label Smoothing (LS) on this benchmark using ResNet-50 as the backbone. Specifically, LS yields a better NLL (1.5730 vs. 1.8431), implying more confident probabilistic predictions. However, MaxSup achieves a better ECE (0.1479 vs. 0.1741), indicating better calibration of the predicted confidence scores. These results validate that MaxSup remains effective on OOD datasets, achieving performance comparable to LS across all three metrics.
Ablation on the Weight Schedule. We also systematically investigate how different $\alpha$ scheduling strategies impact MaxSup’s performance. Empirical results indicate that MaxSup consistently maintains high accuracy across a wide range of schedules, further underscoring its robustness against hyperparameter changes. For additional details and discussions, refer to Appendix˜G.
@@
Label Smoothing 42.4 (+0.3+0.3)
MaxSup 42.8 (+0.7+0.7)
@@ 4.3 Evaluation on Semantic Segmentation
We further investigate MaxSup’s applicability to downstream tasks by evaluating its performance on semantic segmentation using the widely adopted MMSegmentation framework. [^3] https://github.com/open-mmlab/mmsegmentation Specifically, we adopt the UperNet [xiao2018unified] architecture with a DeiT-Small backbone, trained on ADE20K. Models pretrained on ImageNet-1K with either MaxSup or Label Smoothing are then fine-tuned under the same cross-entropy objective (Section˜4.2.2).
@@
Figure 2: Grad-CAM [Selvaraju_2019] visualizations for DeiT-Small models under three training setups: MaxSup (2nd row), Label Smoothing (3rd row), and a baseline (4th row). The first row shows the original images. Compared to Label Smoothing, MaxSup more effectively filters out non-target regions and highlights essential features of the target class, reducing instances where the model partially or entirely focuses on irrelevant areas.
@@ 4.4 Visualization via Class Activation Maps
To better understand how MaxSup fundamentally differs from Label Smoothing (LS) in guiding model decisions, we employ Gradient-weighted Class Activation Mapping (Grad-CAM) [Selvaraju_2019], which highlights regions most influential for each prediction.
We evaluate DeiT-Small under three training setups: MaxSup (second row), LS (third row), and a baseline with standard cross-entropy (fourth row). As illustrated in Figure˜2, MaxSup-trained models more effectively suppress background distractions than LS, which often fixates on unrelated objects—such as poles in “Bird,” tubes in “Goldfish,” and caps in “House Finch.” This behavior reflects LS’s error-enhancement mechanism, which can misdirect attention.
Moreover, MaxSup retains a wider spectrum of salient features, as exemplified in the Shark” and Monkey” images, where LS-trained models often omit crucial semantic details (e.g., fins, tails, or facial contours). These findings align with our analysis in Appendix˜J, clearly demonstrating that MaxSup preserves richer intra-class information. Consequently, MaxSup-trained models produce more accurate and consistent predictions by effectively leveraging fine-grained object cues. Further quantitative Grad-CAM overlay metrics (e.g., precision and recall for target regions) confirm that MaxSup yields more focused and comprehensive activation maps, further underscoring its overall efficacy.
@@ 5 Conclusion
Limitations. Prior work [muller2019does] notes that LS-trained teachers may degrade knowledge distillation [hinton2015distilling, Hu2021CVPR], and guo2024cross suggests LS accelerates convergence via improved conditioning. Examining MaxSup’s potential role in distillation and its overall impact on training dynamics would clarify these underlying effects. Recent studies [sukenik2024neural, garrod2024persistence] also show that $\ell_{2}$ regularization biases final-layer features toward low-rank solutions, raising interesting questions about whether MaxSup behaves similarly.
@@
Acknowledgments
This work was supported by the University of Washington Faculty Startup Fund, the Carwein–Andrews Fellowship, the UW GSFEI Top Scholar Award, and the U.S. DOT PacTrans sub-center seed funding program. We also thank the anonymous reviewers for their constructive feedback and insightful suggestions, which helped improve the quality and clarity of this paper.
@@ 63. Answer: [Yes]
Appendix A Technical Appendices and Supplementary Material
Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline. There is no page limit for the technical appendices.
Appendix B Proof of Lemma 3.2
@@ Appendix A Technical Appendices and Supplementary Material
Appendix C Proof of Theorem˜3.3
@@
Appendix D Gradient Analysis
D.1 New Objective Function
@@ D.2 Gradient Analysis
Appendix E Pseudo Code
@@
Appendix F Robustness Under Different Training Recipes
@@
Appendix G Increasing Smoothing Weight Schedule
@@
As shown in Table˜11, both Label Smoothing and MaxSup benefit from this $\alpha$ scheduler. For Label Smoothing, accuracy improves from 75.91% to 76.16%, while MaxSup sees a more pronounced gain, from 76.12% to 76.58%. This greater improvement for MaxSup ($0.46\%$) compared to Label Smoothing ($0.25\%$) corroborates our claim that MaxSup successfully addresses the inconsistent regularization and error-enhancement issues of Label Smoothing during misclassifications.
@@
Appendix H Extended Evaluation of Linear Transferability on Different Datasets
To further demonstrate the substantial improvement in feature representation compared to other methods, we further compare the linear transfer accuracies of different methods on a broader range of datasets in Table˜12
Table 12: Validation performance of different methods, evaluated using multinomial logistic regression with l2 regularization. Although Label Smoothing and OLS improve ImageNet accuracy, they substantially degrade transfer accuracy compared to MaxSup. Following [kornblith2021better], we selected from 45 logarithmically spaced values between $10^{-6}$ and $10^{5}$. Notably, the search range is larger than the search range used in Table 3, thus leading to higher overall accuracies on CIFAR10.
@@ Appendix H Extended Evaluation of Linear Transferability on Different Datasets
Appendix I Extended Comparison to More Label Smoothing Alternatives
We have included Confidence Penalty [pereyra2017regularizing] and Adaptive Label Smooothing with Self-Knowledge [lee2022adaptive] for an extended comparison in Table˜13. We follow their recommended hyperparameter settings: weight coefficient is set to 0.1 for confidence penalty, and the checkpoint with the highest validation accuracy checkpoint is treated as teacher for Adaptive Label Smoothing.
To further address the novelty of our work, we additionally compared our method to recent approaches that also identify and aim to fix the issues of Label Smoothing on misclassified samples [gao2020towards, wei2022mitigatingneuralnetworkoverconfidence]. Specifically, the method proposed in [wei2022mitigatingneuralnetworkoverconfidence] aims to mitigate overconfidence via logit normalization during training. With its default hyperparameter settings, it achieves an accuracy of 74.32% on ImageNet with a ResNet-50, which is significantly lower than the 76.91% achieved by standard Label Smoothing. This performance aligns with that of Logit Penalty, which similarly minimizes the global $l_{2}$-norm of logits and can struggle to match baseline LS performance. We also note that these norm-based methods[gao2020towards, wei2022mitigatingneuralnetworkoverconfidence, dauphin2021deconstructing] are often highly sensitive to hyperparameter choices, which can limit their practical applicability.
As shown in 13, MaxSup outperforms all these alternatives on ImageNet with ResNet-50. This aligns with our theoretical analysis that selectively penalizing $z_{max}$ yields a more consistent and effective regularization than penalizing all logits (Confidence Penalty), dynamically smoothing the label distribution with self-knowledge (Adaptive LS), or applying global norm-based penalties (Logit Penalty, Logit Normalization).
Table 13: Comparison of classic convolutional neural networks on ImageNet. Results are reported as “mean $\pm$ std” (percentage). Bold entries highlight the best performance; underlined entries mark the second best. (Methods with denote code adaptations from official repositories; see text for details.)